# ROSE: Register-assisted General Time Series Forecasting with Decomposed Frequency Learning

### v2

@@ A.8 Results deviation
ETTm1 0.342$\pm$0.003 0.367$\pm$0.002 0.349$\pm$0.003 0.372$\pm$0.002 0.349$\pm$ 0.004 0.383$\pm$0.003 99%
ETTm2 0.246$\pm$0.002 0.303$\pm$0.004 0.249$\pm$0.002 0.308$\pm$0.002 0.255$\pm$0.002 0.314$\pm$0.003 99%
ETTh1 0.392$\pm$0.004 0.413$\pm$0.004 0.397$\pm$0.003 0.419$\pm$0.003 0.411$\pm$0.003 0.432$\pm$0.005 99%
ETTh2 0.330$\pm$0.003 0.374$\pm$0.002 0.335$\pm$0.004 0.380$\pm$0.003 0.348$\pm$0.004 0.390$\pm$0.004 99%
Traffic 0.391$\pm$0.008 0.266$\pm$0.005 0.418$\pm$0.011 0.278$\pm$0.006 0.404$\pm$0.009 0.283$\pm$0.002 99%
Weather 0.217$\pm$0.008 0.250$\pm$0.007 0.224$\pm$0.007 0.252$\pm$0.009 0.223$\pm$0.011 0.263$\pm$0.014 99%
Electricity 0.156$\pm$0.007 0.249$\pm$0.009 0.164$\pm$0.004 0.253$\pm$0.004 0.163$\pm$0.009 0.261$\pm$0.013 99%
@@ A.9.2 Visualization showcase
To provide a distinct comparison among different models, we present visualizations of the forecasting results on the ETTh2 dataset and the weather dataset in different settings, as shown in Figures 12 to Figures 15, given by the following models: DLinear [Zeng et al., 2023], TimesNet [Wu et al., 2022], iTransfomrer [Liu et al., 2023], and PatchTST [Nie et al., 2022]. Among the methods, ROSE demonstrates the most accurate prediction ability.
@@ A.10.1 Full-shot results
Models ROSE ITransformer PatchTST Timesnet Dlinear GPT4TS S2P-LLM
@@ A.10.2 Few-shot results
Models ROSE ITransformer PatchTST Timesnet Dlinear GPT4TS S2P-LLM
@@ A.10.3 Ablation study results
Novelty of decomposed frequency learning. Frequency masking is not a new concept, but past approaches randomly mask frequencies of a single time series once  [Chen et al., 2023b, Zhang et al., 2023], which show limited forecasting effectiveness due to the lack of common pattern learning from heterogeneous time series that come from various domains. While the multi-frequency masking we proposed randomly mask either high-frequency or low-frequency components of a time series multiple times as the key to enable learning of common time series patterns, such as trends and various long and short term fluctuations. Moreover, different from utilizing frequency masking as a way of data augmentation to enhance the diversity of input data  [Chen et al., 2023b, Zhang et al., 2023], we combine multi-frequency masking with reconstruction task as a novel pre-training framework, that learns a universal and unified feature representation by comprehending the data from various frequency perspectives, thereby enabling it to learn generalized representations.
Difference between frequency-domain masking and time-domain noise addition. Multi-frequency masking and reconstruction are not equivalent to the pre-training methods of adding noise and denoising (Noise). Due to the sparsity of time series, the process of adding noise and denoising may potentially disrupt the information of original time series [Dong et al., 2024]. In contrast, multi-frequency masking not only preserves the series from such disruption but also helps the model understand temporal patterns from a multi-frequency perspective, thereby helping the model to learn general features better.
@@
Replace Multi-Frequency Masking Random Frequency Masking 96 0.330 0.370 0.170 0.262 0.391 0.399 0.303 0.359
@@
Other pre-training tasks Aug 96 0.304 0.357 0.178 0.266 0.376 0.405 0.281 0.348

### v3

@@ A.8 Results Deviation
[…]
@@ A.9.2 Visualization showcase
To provide a distinct comparison among different models, we present visualizations of the forecasting results on the ETTh2 dataset and the weather dataset in different settings, as shown in Figures 12 to Figures 15, given by the following models: DLinear (Zeng et al., 2023), TimesNet (Wu et al., 2022), iTransfomrer (Liu et al., 2023), and PatchTST (Nie et al., 2022). Among the methods, ROSE demonstrates the most accurate prediction ability.
@@ A.10.1 Full-shot results
Table 12 shows the full results of ROSE in full-shot setting for four prediction lengths. ROSE shows the advantage over the specific models and LLM-based models trained with the full training set.
@@
Models ReadyTS ITransformer PatchTST Timesnet Dlinear GPT4TS $\textbf{S}^{2}$IP-LLM
@@ A.10.2 Few-shot results
Table 13 shows the full results of ROSE in 10% few-shot setting for four prediction lengths. ROSE shows the advantage over the specific models and LLM-based models trained with the 10% training set.
@@
Models ReadyTS ITransformer PatchTST Timesnet Dlinear GPT4TS $\textbf{S}^{2}$IP-LLM
@@ A.10.3 Ablation study results
Novelty of decomposed frequency learning. Frequency masking is not a new concept, but past approaches randomly mask frequencies of a single time series once  (Chen et al., 2023b; Zhang et al., 2023), which show limited forecasting effectiveness due to the lack of common pattern learning from heterogeneous time series that come from various domains. While the multi-frequency masking we proposed randomly mask either high-frequency or low-frequency components of a time series multiple times as the key to enable learning of common time series patterns, such as trends and various long and short term fluctuations. Moreover, different from utilizing frequency masking as a way of data augmentation to enhance the diversity of input data  (Chen et al., 2023b; Zhang et al., 2023), we combine multi-frequency masking with reconstruction task as a novel pre-training framework, that learns a universal and unified feature representation by comprehending the data from various frequency perspectives, thereby enabling it to learn generalized representations.
Difference between frequency-domain masking and time-domain noise addition. Multi-frequency masking and reconstruction are not equivalent to the pre-training methods of adding noise and denoising (Noise). Due to the sparsity of time series, the process of adding noise and denoising may potentially disrupt the information of original time series (Dong et al., 2024). In contrast, multi-frequency masking not only preserves the series from such disruption but also helps the model understand temporal patterns from a multi-frequency perspective, thereby helping the model to learn general features better.
@@
Replace
Multi-Frequency Masking
Random Frequency Masking 96 0.330 0.370 0.170 0.262 0.391 0.399 0.303 0.359
@@
Other
pre-training tasks
Aug 96 0.304 0.357 0.178 0.266 0.376 0.405 0.281 0.348
@@ A.10.4 Zero-shot results
Table 15 shows the full results of ROSE and other foundation models in zero-shot setting for four prediction lengths. ROSE exhibits competitive performance.